OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! Solve the system of equations and choose the correct answer from the list of options. d + e = 1 −d + e = −5 Label the ordered pair as (d, e). (0, 0) (3, −2) (−2, 3)

OpenStudy (anonymous):

(−3, 0) is the last answer choice

OpenStudy (anonymous):

ITS (3,-2) 3-2=1 -3+(-2)=-5

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you are welcome